# ToolHaystack: Stress-Testing Tool-Augmented Language Models in Realistic Long-Term Interactions

### v1

@@ A.2 Benchmarks for Evaluating TALMs
As TALMs have advanced, various benchmarks have been introduced to evaluate their capabilities in using external APIs. Early benchmarks primarily assessed models in static, single-turn settings without environmental feedback. For instance, API-Bank (Li et al., 2023) provides 314 task scenarios across 73 APIs, measuring planning, selection, and calling accuracy. Similarly, APIBench (Patil et al., 2023) evaluates whether a model can generate syntactically and semantically correct API calls over a large corpus of public functions. These benchmarks test the model’s symbolic reasoning and API schema understanding but lack interactive components.
Subsequent work introduces real-world interaction into the evaluation loop, enabling models to be assessed in dynamic environments. RestGPT (Song et al., 2023) connects LLM agents to actual RESTful APIs and evaluates their ability to plan API sequences and recover from execution errors. To further approximate user-facing scenarios, recent benchmarks have introduced user simulators and multi-turn dialogues. ToolDial (Shim et al., 2025) includes over 11k dialogues annotated with 16 action types, capturing realistic tool-assisted assistant behavior with clarifications, failures, and API chaining. $\tau$-bench (Yao et al., 2025) extends this idea by simulating domain-constrained interactions (e.g., shopping, airline booking) and measuring goal completion in a transactional database, along with a consistency metric (pass@$k$) to evaluate reliability across repeated trials.
Recent benchmarks have introduced more complex setups to test advanced capabilities like planning and tool use. UltraTool (Huang et al., 2024) separates planning from execution to assess models’ ability to devise tool-use plans. ToolSandbox (Lu et al., 2024) adds stateful tools and on-policy user simulators, supporting persistent states and multi-tool chaining. HammerBench (Wang et al., 2024) expands on this with diverse multi-turn tasks, including Q&A and argument shifts. ComplexFuncBench (Zhong et al., 2025) addresses the challenge of extremely long contexts (e.g., 128k tokens) and proposes constructing long function call sequences.
@@ C.5 Interaction Simulation
C.6 Quality Assurance
We employ multiple frontier LLMs (GPT-4o, Claude 3.5, Gemini 2.5, LLaMA 3.3) during generation to reduce model-specific artifacts. Automated validation enforces constraints such as correct argument use, proper tool syntax, and scenario fidelity. A subset of generated dialogues undergoes a filtering process, where we prompt gpt-4o to strictly evaluate if the dialogue violates the rules for the target category.
Moreover, to ensure the quality of ToolHaystack, we employed GPT-4o as a judge to automatically filter out dialogues that do not satisfy our scenario definitions. To validate the effectiveness of this automated filtering process, we sampled 60 dialogues (10 from each category) from the unfiltered dataset and compared the results against human annotations. The evaluation yielded a Matthews Correlation Coefficient (MCC) of 0.884, indicating a strong agreement between the automated and human filtering outcomes.
C.7 Dataset Generation Details
For ContextRecall-$multi$ scenario, we find that if APIs used for needle include multiple parameters that work similarly with any parameter in the target API, it can repeatedly appear in the dialogue context and makes it easier to find the needle. To this reason, we utilize API triplets where APIs used for needle have a single shared argument with the target API.
@@ Appendix D Evaluation Details
We use greedy decoding (i.e., temperature set to 0) and limit the maximum number of generated tokens to 2,000. For closed-source models, we access the GPT family via the OpenAI API and the Claude models via the Anthropic API. Other models accessible through OpenRouter are used via that service. Finetuned TALMs are served using vLLM (Kwon et al., 2023).
@@
D.1 Case Study
Table 8 shows an example failure of GPT-4o-mini.
Sample Dialogue from Missing Context-Hard:
(Previous sessions)
Ground-truth Response:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "MISSING"
GPT-4o-mini Prediction:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "3"
DeepSeek-V3 Output:
  "api_name": "Search domains",
  "parameters": {
    "q": "suspiciousdomain.com",
    "type": "WEBIP",
    "pagenum": "1"

### v2

@@ A.2 Benchmarks for Evaluating TALMs
As TALMs have advanced, various benchmarks have been introduced to evaluate their capabilities in using external APIs. Early benchmarks primarily assessed models in static, single-turn settings without environmental feedback. For instance, API-Bank (li2023apibank) provides 314 task scenarios across 73 APIs, measuring planning, selection, and calling accuracy. Similarly, APIBench (patil2023gorilla) evaluates whether a model can generate syntactically and semantically correct API calls over a large corpus of public functions. These benchmarks test the model’s symbolic reasoning and API schema understanding but lack interactive components.
Subsequent work introduces real-world interaction into the evaluation loop, enabling models to be assessed in dynamic environments. RestGPT (song2023restgpt) connects LLM agents to actual RESTful APIs and evaluates their ability to plan API sequences and recover from execution errors. To further approximate user-facing scenarios, recent benchmarks have introduced user simulators and multi-turn dialogues. ToolDial (shimtooldial) includes over 11k dialogues annotated with 16 action types, capturing realistic tool-assisted assistant behavior with clarifications, failures, and API chaining. $\tau$-bench (yao2025taubench) extends this idea by simulating domain-constrained interactions (e.g., shopping, airline booking) and measuring goal completion in a transactional database, along with a consistency metric (pass@$k$) to evaluate reliability across repeated trials.
Recent benchmarks have introduced more complex setups to test advanced capabilities like planning and tool use. UltraTool (ultratool) separates planning from execution to assess models’ ability to devise tool-use plans. ToolSandbox (lu2024toolsandbox) adds stateful tools and on-policy user simulators, supporting persistent states and multi-tool chaining. HammerBench (wang2024hammerbench) expands on this with diverse multi-turn tasks, including Q&A and argument shifts. ComplexFuncBench (zhong2025complexfuncbench) addresses the challenge of extremely long contexts (e.g., 128k tokens) and proposes constructing long function call sequences.
@@ C.5 Interaction Simulation
C.6 Dataset Generation Details
For ContextRecall-multimulti scenario, we find that if APIs used for needle include multiple parameters that work similarly with any parameter in the target API, it can repeatedly appear in the dialogue context and makes it easier to find the needle. To this reason, we utilize API triplets where APIs used for needle have a single shared argument with the target API.
@@ Appendix D Evaluation Details
We use greedy decoding (i.e., temperature set to 0) and limit the maximum number of generated tokens to 2,000. For closed-source models, we access the GPT family via the OpenAI API and the Claude models via the Anthropic API. Other models accessible through OpenRouter are used via that service. Finetuned TALMs are served using vLLM (kwon2023vllm).
@@
Model
Communication
& Social
Finance
& Business
Lifestyle
& Services
Media
& Entertainment
Utility
& Tools
Other
GPT-4.5-Preview 0.6667 0.3731 0.3077 0.5465 0.4105 0.4384
Claude-3-Opus 0.2051 0.1493 0.0769 0.1279 0.1579 0.1370
DeepSeek-V3 0.2564 0.2239 0.2308 0.3837 0.2421 0.3562
Table 9: Domain-wise evaluation of model performance.
Appendix E Additional Results
We also conducted a domain-wise evaluation of model performance using a taxonomy that clusters fine-grained API categories into six broader domains. We grouped over 30 fine-grained API categories provided by Rapid API into six high-level domains based on their functional similarity. We report the accuracy of three representative models across the six major domains in the Table 9. These studies demonstrate that ToolHaystack enables controlled, interpretable error analysis across multiple dimensions.